# ReorderBench: A Benchmark for Matrix Reordering

### v1

@@ IV-A2 Variation Generation
Adding noise. Behrisch et al.  [21] have found that combining noise in the form of anti-patterns with visual patterns increases the diversity of matrices. Therefore, we utilize such noise to degenerate matrix templates into a set of variations. As shown in Fig. 5, two main intrinsic anti-patterns exist in the matrices: noise and noise-cluster. Noise anti-patterns are characterized by the random distribution of non-zero entries throughout the matrix, which lacks inherent visual patterns. Conversely, noise-cluster anti-patterns form artificial clusters when noise aggregates. Each cluster consists of similar yet not identical rows. Three steps for adding noises are: 1. determining noise levels, 2. noise generation, and 3. noise application.
@@
The upper bound for the noise levels is set to the maximum noise level that can visually preserve the patterns. Through analysis of matrices subjected to noises of various levels, we find that $16\%$ is an appropriate upper bound. The detailed analysis is provided in the supplemental material. Uniformly distributed noise levels are commonly used in dataset generation [33]. Therefore, we use a set of uniform noise levels: $[0\%,1\%,\dots,16\%]$. For each matrix variation, we randomly select two noise levels from this set, one for each type of anti-pattern. For each matrix template, a sufficient number of variations should be generated to cover the set of noise levels. We derive in the supplemental material that approximately $70$ variations are required in our case.
Based on the selected noise levels, we generate noise anti-patterns and noise-cluster anti-patterns separately. For noise anti-patterns, we randomly introduce non-zero entries throughout the matrix according to the noise level. For noise-cluster anti-patterns, a straightforward method is to generate several clusters of similar yet not identical rows and then apply index swaps to simultaneously distribute the cluster-associated rows throughout the matrix. However, in such anti-patterns, all artificial clusters can be recovered simultaneously by reversing the index swaps. They will severely distort the visual patterns in the template. To tackle this issue, we generate and distribute each cluster separately so that they will be recovered by different orderings of the matrix rows and columns. Specifically, as shown in Fig. 6, we first generate a set of $n$-dimensional noise vectors, where $n$ is the number of rows in the matrices. Each noise vector randomly contains several non-zero entries determined by the specified noise level. Then, for each row of the anti-pattern, we randomly select one of these noise vectors. The rows with the same noise vector form a cluster. Finally, we symmetrize the anti-pattern by performing a logical OR with its transposed version. This symmetrization process ensures that rows within the same cluster are similar yet not identical. To moderate the impact of noise-cluster anti-patterns on visual patterns across matrix templates, choosing an appropriate set size for the noise vectors is crucial. If the set is too small, a single noise vector can appear in too many rows. As a result, the rows with this noise vector form large clusters, which severely distort the visual patterns in the template. On the other hand, if the set is too large, each noise vector appears in too few rows, making the noise-cluster anti-pattern nearly identical to noise anti-patterns. After carefully studying $12,000$ generated matrices with anti-patterns, we have observed that the set size approximately equals the average number of rows and columns across all visual patterns.
@@
Swapping indices. To further increase the diversity of the matrix variations, we randomly swap their rows and columns to simulate real-world matrices that are not fully reordered. The upper bound for the number of index swaps should be sufficient to make the order of matrix rows and columns completely random. In the supplemental material, we derive that this upper bound is $\frac{1}{2}n\log n$, where $n$ is the number of rows in the matrix. As the number of index swaps increases, the rate of degeneration slows down, leading to a more gradual decline in the visual pattern quality. Therefore, we use numbers that increase exponentially rather than linearly to ensure different degrees of quality in visual patterns. Specifically, for each matrix variation, we apply varying numbers of index swaps, including 00 and powers of 2 up to the closest one to $\frac{1}{2}n\log n$.
@@ IV-B Score Calculation
A straightforward solution to score the visual patterns in a matrix is to measure the entry-wise similarity with its template. However, this method ignores the translation invariance of visual patterns. For example, translating a block pattern along the main diagonal does not affect its visual quality. However, it can significantly change the entry-wise similarity. To tackle this issue, we consider how well the regions in the matrix can match a visual pattern and how well they present the visual pattern as a connected component. Previous studies have shown that 1. convolutional kernels are capable of matching visual features in images [6], and 2. entropy effectively quantifies the level of disorder in a set of potential outputs [7]. Building on these findings, we develop a convolution- and entropy-based scoring method that combines the matching capability of convolution and the disorder detection capability of entropy. This scoring method consists of two phases: pattern detection and pattern scoring. In the pattern detection phase, given a matrix variation and a visual pattern from its template, we utilize the matching capability of convolutional kernels to greedily match this pattern in the matrix variation and identify the region with the highest matching score (Fig. 7). In the pattern scoring phase, we adjust this matching score based on the connectedness of the matched region, an important aspect of its visual quality [34]. The disorder detection capability of entropy is effective in quantifying this connectedness. Therefore, we use it to calculate a disorder score for the matched region. For continuous matrices, we further include the deviation score to measure the deviation of the matched region from the Robinson structure. The final score of a matrix variation is derived from the scores of all matched regions.
Pattern detection. The pattern detection phase detects visual patterns in both binary and continuous matrices. To facilitate this process, continuous matrices are converted to binary ones by setting non-zero entries to $1$. As shown in Fig. 8, the convolutional kernel for each pattern is first derived as a matrix that contains only the pattern itself. In particular, for off-diagonal block and band patterns that consist of diagonal symmetric components, we simplify the kernel to contain only one of the components. The convolutional kernel is then used to scan the matrix variation and calculate convolutions with its regions, as illustrated in Fig. 7. For block and star patterns, the kernel scans along the main diagonal. For off-diagonal block and band patterns, the kernel scans the off-diagonal regions. The region with the highest convolution that does not overlap with previously matched regions is chosen to match the pattern. Since larger visual patterns reveal more meaningful information, the matching process starts with the largest patterns and proceeds to the smallest ones.
Figure 7: The greedy matching strategy in the pattern detection phase. A convolutional kernel scans the matrix variation and the region with the highest matching score is chosen.
@@
Pattern scoring. In the pattern scoring phase, the matching scores, disorder scores, and deviation scores are derived to measure the quality of visual patterns presented by the matched regions. These scores are then aggregated to produce a final score for the matrix variation.
The matching score ($S^{m}$) is measured by the convolution between the matched region and the convolutional kernel. To make the scores comparable across visual patterns, this convolution is normalized by the area of the matched region. Formally, for a matched region with entries $\bm{a}=\{a_{i,j}\}$, the matching score is calculated as $S^{m}=\sum_{i,j}\mathbb{I}[a_{i,j}>0]/|\bm{a}|$, where $\mathbb{I}$ is the indicator function that takes the value 1 if its argument is true and 0 otherwise, and $|\bm{a}|$ is the number of entries in the matched region, representing its area.
@@
The disorder score ($S^{d}$) measures the degree to which the matched region is fragmented into smaller components by zero entries. As shown in Fig. 9, this score is derived from the entropy of the proportions of the connected non-zero components. To make the scores comparable across visual patterns, we normalize the entropy by the logarithm of the area of the matched region. Formally, for a matched region with $\ell$ connected components each occupying $pr_{i}$ percentage of the total non-zero entries, the disorder score is calculated as $S^{d}=(-\sum_{i=1}^{\ell}pr_{i}\cdot\log(pr_{i}))/\log(|\bm{a}|)$.
@@
The deviation score ($S^{v}$) quantifies the degree to which entries violate the Robinson condition of monotonically decreasing away from the diagonal (Fig. 10). To make the scores comparable across visual patterns, we normalize the deviation score by the maximum possible deviation. For block and star patterns, it is calculated as:
@@
Here, $i<k<j$. $\mathbb{I}[a_{i,k}<a_{i,j}]$ and $\mathbb{I}[a_{k,j}<a_{i,j}]$ indicates whether two entries violate the Robinson condition, while $f(a_{i,k},a_{i,j})$ and $f(a_{k,j},a_{i,j})$ measures the degree of violation. Specifically, $f(a_{i,k},a_{i,j})=\mathbb{I}[a_{i,k}>0]\mathbb{I}[a_{i,j}>0]|a_{i,k}-a_{i,j}|$ ensures that the violation is counted only when both entries are non-zero, as the disorder score already heavily penalizes the zero entries for fragmenting the components. For off-diagonal block patterns with multiple diagonals, we identify the diagonal that minimizes the deviation score. Therefore, we calculate the score as the minimum of $S^{v}$ values derived from all diagonals. Please refer to the supplemental material for more details. For binary matrix variations, we set the deviation scores to 0.
A matched region that simultaneously achieves a high matching score, a low disorder score, and a low deviation score presents a high-quality visual pattern. Thus, the quality score ($S$) of a matched region is calculated as:
$S=S^{m}\cdot(1-S^{d})\cdot(1-S^{v}).$ (2)
@@ IV-C Statistics
We have generated the largest benchmark, ReorderBench, for matrix reordering tasks. It contains $8,505,000$ matrices, including $2,835,000$ binary and $5,670,000$ continuous matrices. Each matrix comes with a score to reflect the quality of visual patterns. These matrices are of four sizes: [$100{\mkern-2.0mu\times\mkern-2.0mu}100$, $200{\mkern-2.0mu\times\mkern-2.0mu}200$, $300{\mkern-2.0mu\times\mkern-2.0mu}300$, $400{\mkern-2.0mu\times\mkern-2.0mu}400$]. The upper bound for the matrix size is set to $400{\mkern-2.0mu\times\mkern-2.0mu}400$ because it approaches the maximum size that can be effectively displayed on standard screens. For instance, on a display with a resolution of 1920×1080, each entry in such matrices can be only $\lfloor 1080/400\rfloor=2$ pixels wide. Fig. 11 shows examples of generated matrices. For more examples, please refer to the supplemental material and our website at https://reorderbench.github.io/.
In ReorderBench, the matrix distribution across different attributes, such as noise level and size, is relatively balanced. For example, variations in the percentage of matrices at each noise level do not exceed $0.05\%$. Fig. 12 shows that the generated matrices effectively cover different score ranges of visual patterns. While star and band patterns tend to have more low-score matrices due to their susceptibility to noise and index swaps, the overall distribution of matrices across pattern types and scores remains balanced, with the largest difference being $629,069:226,372\approx 2.78:1$. This is noticeably lower than the commonly referenced $4:1$ standard [35].
@@
We split ReorderBench into a training set of $6,804,000$ matrices and a test set of $1,701,000$ matrices while maintaining the same distribution across attributes in both sets. To ensure that reordering methods are evaluated on matrix templates unseen during training, we assign each matrix variation to the same set as its corresponding template.
To accommodate tasks that require different distributions, adjustments can be made by sampling or generating more matrices for specific types. Overall, our benchmark offers a comprehensive resource catering to the needs of both researchers and practitioners.
@@ V Data Study on Scoring Consistency
Consistency with human assessments of visual pattern quality greatly influences the usefulness of the scoring method. Therefore, we evaluate how well our convolution- and entropy-based method align with expert assessments. In alignment with prior research [36], we conduct an empirical data study where a few selected experts examine many data, in contrast to the traditional method where a large group of ordinary people reviews limited data. We invite three experts, each with extensive experience in matrix reordering. The first expert is a senior researcher at a major IT company whose research interests include computer vision and visual analytics. He often uses matrix visualization to analyze the log data produced by machine learning models. The other two experts are fifth-year Ph.D. students majoring in visual analytics who are not co-authors of this paper. They have used various reordering algorithms in their research projects.
@@ V-A Study Settings
Baseline. We compare our scoring method against existing quality metrics, which use either the adjacency matrix or the dissimilarity matrix to measure visual pattern quality [25]. Since there is no consensus on their comparative effectiveness, we select two commonly used quality metrics from each category for our analysis. From adjacency-based metrics, we select BW [24] and ME [23]. From dissimilarity-based metrics, we select AR events [27] and BAR [2].
Task and experiment design. To gather the assessments of human experts on visual pattern quality, the task is to compare different variations generated from a given matrix. In the pilot study, for each visual pattern, we sample three binary matrices and three continuous matrices without index swaps, resulting in a total of 24 matrices. For each matrix, we generate eight different variations and evaluate them using both baseline metrics and our method. Four of the variations are generated by applying 00, $8$, $64$, and $512$ index swaps, respectively. The other four variations are generated by optimizing the four baseline metrics. We create questions for all pairwise comparisons of the eight variations, resulting in 28 questions per matrix. Thus, we have $0163(experts)\times 24(matrices)\times 28(questions)=2,$ 016 results from the pilot study. In the formal study, we sample $48$ matrices evenly distributed across patterns, leading to $4,032$ results in the formal study. The matrices involved in the data study are provided in the supplemental material. Based on the gathered assessments, we compare the level of consistency between each quality metric and human experts.

### v2

@@ IV-A2 Variation Generation
Adding noise. Behrisch et al.  [23] have found that combining noise in the form of anti-patterns with visual patterns increases the diversity of matrices. Therefore, we utilize such noise to degenerate matrix templates into a set of variations. As shown in Fig. 5, two main intrinsic anti-patterns exist in the matrices: noise and noise-cluster. Noise anti-patterns are characterized by the random distribution of non-zero entries throughout the matrix, which lacks inherent visual patterns. Conversely, noise-cluster anti-patterns form artificial clusters when noise aggregates. Each cluster consists of similar yet not identical rows. The key distinction between the noise-cluster anti-patterns and actual clusters in the matrix template lies in their purpose in the matrix generation process. Actual clusters are intended to convey meaningful relationships, whereas noise-cluster anti-patterns serve to degenerate patterns in the template by simulating structured noise or anomalies within the data. Addressing noise-cluster anti-patterns is crucial for improving reordering performance, as they frequently obscure meaningful patterns and challenge existing methods [23]. Three steps for adding noises are: 1. determining noise levels, 2. noise generation, and 3. noise application.
@@
The upper bound for the noise levels is set to the maximum noise level that can visually preserve the patterns. Through analysis of matrices subjected to noises of various levels, we find that $16\%$ is an appropriate upper bound. The detailed analysis is provided in Sec. 2 of the supplemental material. Uniformly distributed noise levels are commonly used in dataset generation [40]. Therefore, we use a set of uniform noise levels: $[0\%,1\%,\dots,16\%]$. For each matrix variation, we randomly select two noise levels from this set, one for each type of anti-pattern. For each matrix template, a sufficient number of variations should be generated to cover the set of noise levels. We derive in Sec. 3 of the supplemental material that approximately $70$ variations are required in our case.
Based on the selected noise levels, we generate noise anti-patterns and noise-cluster anti-patterns separately. For noise anti-patterns, we randomly introduce non-zero entries throughout the matrix according to the noise level. For noise-cluster anti-patterns, a straightforward method is to generate several clusters of similar yet not identical rows and then apply index swaps to simultaneously distribute the cluster-associated rows throughout the matrix. However, in such anti-patterns, all artificial clusters can be recovered simultaneously by reversing the index swaps. They will severely distort the visual patterns in the template. To tackle this issue, we generate and distribute each cluster separately so that they will be recovered by different orderings of the matrix rows and columns. Specifically, as shown in Fig. 6, we first generate a set of $n$-dimensional noise vectors, where $n$ is the number of rows in the matrices. Each noise vector randomly contains several non-zero entries determined by the specified noise level. Then, for each row of the anti-pattern, we randomly select one of these noise vectors. The rows with the same noise vector form a cluster. Finally, we symmetrize the anti-pattern by performing a logical OR with its transposed version. This symmetrization process ensures that rows within the same cluster are similar yet not identical. To moderate the impact of noise-cluster anti-patterns on visual patterns across matrix templates and ensure they do not unintentionally form meaningful visual patterns, choosing an appropriate set size for the noise vectors is crucial. If the set is too small, a single noise vector can appear in too many rows. As a result, the rows with this noise vector form large clusters, which severely distort the visual patterns in the template. On the other hand, if the set is too large, each noise vector appears in too few rows, making the noise-cluster anti-pattern nearly identical to noise anti-patterns. After carefully studying $12,000$ generated matrices with anti-patterns, we have observed that the set size approximately equals the average number of rows and columns across all visual patterns. This appropriate set size ensures that noise-cluster anti-patterns remain distinct from both meaningful visual patterns and random noise.
@@
Swapping indices. To further increase the diversity of the matrix variations, we randomly swap their rows and columns to simulate real-world matrices that are not fully reordered. The upper bound for the number of index swaps should be sufficient to make the order of matrix rows and columns completely random. In Sec. 4 of the supplemental material, we derive that this upper bound is $\frac{1}{2}n\log n$, where $n$ is the number of rows in the matrix. As the number of index swaps increases, the rate of degeneration slows down, leading to a more gradual decline in the visual pattern quality. Therefore, we use numbers that increase exponentially rather than linearly to ensure different degrees of quality in visual patterns. Specifically, for each matrix variation, we apply varying numbers of index swaps, including 00 and powers of 2 up to the closest one to $\frac{1}{2}n\log n$.
@@ IV-B Score Calculation
A straightforward solution to score the visual patterns in a matrix is to measure the entry-wise similarity with its template. However, this method ignores the translation invariance of visual patterns. For example, translating a block pattern along the main diagonal does not affect its visual quality. However, it can significantly change the entry-wise similarity. To tackle this issue, we consider how well the regions in the matrix can match a visual pattern and how well they present the visual pattern as a connected component. Previous studies have shown that 1. convolutional kernels are capable of matching visual features in images [6], and 2. entropy effectively quantifies the level of disorder in a set of potential outputs [7]. Building on these findings, we develop a convolution- and entropy-based scoring method that combines the matching capability of convolution and the disorder detection capability of entropy. This scoring method consists of two phases: pattern detection and pattern scoring. In the pattern-detection phase, given a matrix variation and a visual pattern from its template, we utilize the matching capability of convolutional kernels to greedily match this pattern in the matrix variation and identify the region with the highest existence score (Fig. 7). In the pattern-scoring phase, we adjust this existence score based on the connectedness of the matched region, an important aspect of its visual quality [41]. The disorder detection capability of entropy is effective in quantifying this connectedness. Therefore, we use it to calculate a disorder score for the matched region. For continuous matrices, we further include the deviation score to measure the deviation of the matched region from the Robinson structure. The final score of a matrix variation is derived from the scores of all matched regions.
Figure 7: The greedy matching strategy in the pattern-detection phase. A convolutional kernel scans the matrix variation and the region with the highest existence score is chosen.
@@
Pattern detection. The pattern-detection phase detects visual patterns in both binary and continuous matrices. To facilitate this process, continuous matrices are converted to binary ones by setting non-zero entries to $1$, ensuring the focus remains on the existence of meaningful values rather than their exact magnitudes. Fig. 8 illustrates the process of deriving convolutional kernels from each pattern in the matrix template. A convolutional kernel is represented as a matrix containing only the pattern itself. In particular, for off-diagonal block and band patterns that consist of diagonal symmetric components, we simplify the kernel to contain only one of the components. The convolutional kernel is then used to scan the matrix variation and calculate convolutions with its regions, as illustrated in Fig. 7. For block and star patterns, the kernel scans along the main diagonal. For off-diagonal block and band patterns, the kernel scans the off-diagonal regions. The region with the highest convolution that does not overlap with previously matched regions is chosen to match the pattern. Since larger visual patterns reveal more meaningful information, the matching process starts with the largest patterns and proceeds to the smallest ones.
Pattern scoring. In the pattern-scoring phase, the existence scores, disorder scores, and deviation scores are derived to measure the quality of visual patterns presented by the matched regions. These scores are then aggregated to produce a final score for the matrix variation.
The existence score ($S^{e}$) is measured by the convolution between the matched region and the convolutional kernel. To make the scores comparable across visual patterns, this convolution is normalized by the area of the matched region. Formally, for a matched region with entries $\bm{a}=\{a_{i,j}\}$, the existence score is calculated as ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}S^{e}}=\sum_{i,j}\mathbb{I}[a_{i,j}>0]/|\bm{a}|$, where $\mathbb{I}$ is the indicator function that takes the value 1 if its argument is true and 0 otherwise, and $|\bm{a}|$ is the number of entries in the matched region, representing its area. Rounding non-zero entries to 1 in this score ensures the focus remains on the existence of meaningful values rather than their exact magnitudes.
@@
The disorder score ($S^{d}$) measures the degree to which the matched region is fragmented into smaller components by zero entries. As shown in Fig. 9, this score is derived from the entropy of the proportions of the connected non-zero components. In computing these components, we consider two entries connected if they share an edge or a corner. To make the scores comparable across visual patterns, we normalize the entropy by the logarithm of the area of the matched region. Formally, for a matched region with $\ell$ connected components each occupying $pr_{i}$ percentage of the total non-zero entries, the disorder score is calculated as $S^{d}=(-\sum_{i=1}^{\ell}pr_{i}\cdot\log(pr_{i}))/\log(|\bm{a}|)$.
@@
The deviation score ($S^{v}$) is adapted from AR deviations [28] to evaluate all four types of visual patterns. It quantifies the degree to which entries violate the Robinson condition of monotonically decreasing away from the diagonal (Fig. 10). To make the scores comparable across patterns, we normalize the deviation score by the maximum possible deviation. For block and star patterns, it is calculated as:
@@
Here, $i<k<j$. $\mathbb{I}[a_{i,k}<a_{i,j}]$ and $\mathbb{I}[a_{k,j}<a_{i,j}]$ indicate whether two entries violate the Robinson condition, while $f(a_{i,k},a_{i,j})$ and $f(a_{k,j},a_{i,j})$ measure the degree of violation. Specifically, $f(a_{i,k},a_{i,j})=\mathbb{I}[a_{i,k}>0]\mathbb{I}[a_{i,j}>0]|a_{i,k}-a_{i,j}|$ ensures that the violation is counted only when both entries are non-zero, as the disorder score already heavily penalizes the zero entries for fragmenting the components. For off-diagonal block patterns, there are multiple diagonals along which the Robinson structure can be formed. To address this ambiguity, we identify the diagonal that minimizes the deviation score, which most accurately reflects the underlying Robinson structure [28]. Therefore, we calculate the score as the minimum of $S^{v}$ values derived from all diagonals. Please refer to Sec. 5 of the the supplemental material for more details. For binary matrix variations, we set the deviation scores to 0.
A matched region that simultaneously achieves a high existence score, a low disorder score, and a low deviation score presents a high-quality visual pattern. Thus, the quality score ($S$) of a matched region is calculated as:
$S={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}S^{e}}\cdot(1-S^{d})\cdot(1-S^{v}).$ (2)
@@ IV-C Statistics
We have generated the largest benchmark, ReorderBench, for matrix-reordering tasks. It contains $8,505,000$ matrices, including $2,835,000$ binary and $5,670,000$ continuous matrices. Each matrix comes with a score to reflect the quality of visual patterns. These matrices are of four sizes: [$100{\mkern-2.0mu\times\mkern-2.0mu}100$, $200{\mkern-2.0mu\times\mkern-2.0mu}200$, $300{\mkern-2.0mu\times\mkern-2.0mu}300$, $400{\mkern-2.0mu\times\mkern-2.0mu}400$]. The upper bound for the matrix size is set to $400{\mkern-2.0mu\times\mkern-2.0mu}400$ because it approaches the maximum size that can be effectively displayed on standard screens. For instance, on a display with a resolution of 1920×1080, each entry in such matrices can be only $\lfloor 1080/400\rfloor=2$ pixels wide. Fig. 11 shows examples of generated matrices. For more examples, please refer to Sec. 6 of the the supplemental material and our website at https://reorderbench.github.io/.
@@
[…]
To ensure consistent and reproducible evaluation on ReorderBench, we split it into a training set of $6,804,000$ matrices and a test set of $1,701,000$ matrices while maintaining the same distribution across attributes in both sets. To ensure that reordering methods are evaluated on matrix templates unseen during training, we assign each matrix variation to the same set as its corresponding template.
To accommodate tasks that require different distributions, adjustments can be made by sampling or generating more matrices for specific types. For example, in Sec. 8 of the supplemental material, we demonstrate how our methods can be adapted to overlapping block patterns. To complement the generated matrices, we include $450$ real-world matrices from the Network Repository [12], the TUDataset [44], and the LRGB Dataset [45] in ReorderBench and annotate their visual patterns. Please refer to Sec. 9 of the supplemental material for more details. Overall, our benchmark offers a comprehensive resource catering to the needs of both researchers and practitioners.
@@ V Data Study on Scoring Consistency
Consistency with human assessments of visual pattern quality greatly influences the usefulness of the scoring method. Therefore, we evaluate how well our convolution- and entropy-based method align with expert assessments. In alignment with prior research [46], we conduct an empirical data study where a few selected experts examine many data, in contrast to the traditional method where a large group of ordinary people reviews limited data. We invite three experts, each with extensive experience in matrix reordering. The first expert is a senior researcher at a major IT company whose research interests include computer vision and visual analytics. He often uses matrix visualization to analyze the log data produced by machine learning models. The other two experts are fifth-year Ph.D. students majoring in visual analytics who are not co-authors of this paper. They have used various reordering algorithms in their research projects. The second and third experts collaborate occasionally. There is no in-depth collaboration between the authors and any of the experts. We conduct the data study on both generated and real-world matrices. The study settings and results for the generated matrices are presented here, while those for the real-world matrices are provided in Sec. 9.3 of the supplemental material.
@@ V-A Study Settings
Baseline. We compare our scoring method against existing quality metrics, which either measure the similarity between adjacent entries in the matrix or focus on the distances between the rows and columns in the ordering [21]. Since there is no consensus on their comparative effectiveness, we select commonly used quality metrics for our analysis, including Moran’s I [21], LA [9], AR events [28] and BAR [2].
Task and experiment design. To gather the assessments of human experts on visual pattern quality, the task is to compare different variations generated from a given matrix. In the pilot study, for each visual pattern, we sample three binary matrices and three continuous matrices without index swaps, resulting in a total of 24 matrices. For each matrix, we generate eight different variations and evaluate them using both baseline metrics and our method. Four of the variations are generated by applying 00, $8$, $64$, and $512$ index swaps, respectively. The other four variations are generated by optimizing the four baseline metrics using a simulated annealing-based algorithm, ARSA [15], which exclusively optimizes each metric to ensure the focus on comparing them. We create questions for all pairwise comparisons of the eight variations, resulting in 28 questions per matrix. Thus, we have $0163(experts)\times 24(matrices)\times 28(questions)=2,$ 016 results from the pilot study. In the formal study, we sample $48$ matrices evenly distributed across patterns, leading to $4,032$ results in the formal study. The matrices involved in the data study are provided in Figs. 7, 8, and 9 of the supplemental material. Based on the gathered assessments, we compare the level of consistency between each quality metric and human experts.